Can someone help solve this word problem?
P(x)=2.25x-7,000
a.)What is profit if 2,000 units are sold.
b.)What is profit if 5,000 units are sold.
c.)What is the break even points for sale?
Thank you.

Answer by ankor@dixie-net.com(22740)

P(x)=2.25x-7,000
a.)What is profit if 2,000 units are sold.
Substituting 2000 for x in the given equation
P(x) = 2.25(2000) - 7000
P(x) = 4500 - 7000
P(x) = -2500; not a profit at all, lost $2500
;
b.)What is profit if 5,000 units are sold.
P(x) = 2.25(5000) - 7000
P(x) = 11250 - 7000
P(x) = $4250 profit
:
c.)What is the break even points for sale?
Break-even point is when profit (or loss) = 0
2.25x - 7000 = 0
2.25x = +7000
x = 7000/2.25
x = 3111 units is about the break-even point
:
Not that hard to understand, was it?